OpenStudy (anonymous):

Find the standard from of x^2 + 4y^2 + 8x - 8y + 16 = 0 then find the foci and the center.

OpenStudy (anonymous):

@RadEn can you help??

OpenStudy (raden):

x^2 + 4y^2 + 8x - 8y + 16 = 0 are u sure the second term (4y^2) ? it's not a circle equation

OpenStudy (anonymous):

yes it says 4y^2 and yea it should be and ellipse

OpenStudy (anonymous):

@RadEn are you there?

OpenStudy (raden):

ohhh... sorry i think it's a circle :) x^2 + 4y^2 + 8x - 8y + 16 = 0 x^2 + 8x + 16 + 4y^2-8y = 0 (x+4)^2 + 4(y-1)^2 - 4 = 0 (x+4)^2 + 4(y-1)^2 = 4 divided by 4 to both sides, gives ((x+4)^2)/4 + (y-1)^2 = 1 ((x-(-4)^2)/2^2 + (y-1)^2/(1)^2 = 1 so, the center is (4,1) the foci (c^2) = 2^2 - 1^1 = 4 - 1 = 3

OpenStudy (anonymous):

ok thanks i'll find the major axis and minor by my self

OpenStudy (raden):

yw...yes, right u are

OpenStudy (anonymous):

excuse me but will the foci be 4 and 3?

OpenStudy (raden):

why be 4?

OpenStudy (anonymous):

so it will be -5.73, -2.27 right?

OpenStudy (raden):

how come did u get it, show to m

OpenStudy (anonymous):

that what is says in there

OpenStudy (raden):

lol, wolfram is the best calculator... btw, the typo from me above. the center should be (-4,1) sorry, my bad :)

OpenStudy (anonymous):

okay thanks anyways :)

OpenStudy (raden):

welcome, nice pic^^

OpenStudy (anonymous):

thx